OpenStudy (anonymous):

1. 3x + 3 = 18 2. 2-7x = -47 3. (x/4) + 5 = 7 4. -1 - (2/7) x = -3 5. (x/5) -8 = -6 6. (x/6) + 2 = 2.5

OpenStudy (anonymous):

1. 3x + 3 = 18 x = ( 18-3)/3 x =5

OpenStudy (anonymous):

2. 2-7x = -47 x=(-47-2): -7 x= 7

OpenStudy (anonymous):

4. -1 - (2/7) x = -3 x =(-3+1)*-(7/2) x= 7

OpenStudy (mathteacher1729):

Jones, how far did you get on, say, problem # 3 before getting stuck?

OpenStudy (anonymous):

I'm not sure I see the point in giving the answer to every problem. I don't know math but shouldn't the answer to one be instructive on how to answer the others on his/her own?

OpenStudy (anonymous):

3. (x/4) = 7-5 x/4 = 2 x = 8

OpenStudy (anonymous):

5. (x/5) = -6+8 x/5 = 2 x = 10

OpenStudy (mathteacher1729):

Hi Dgo, I know a little bit of math and you're absolutely correct. :) I hope that Jones is willing to chat a bit and explain where s/he is getting stuck with problems like these so that someone can best help him/her move past that point.